OpenStudy (anonymous):

Write an equation for a parabola with vertex (-6,2) that is congruent to y=1/2x^2 and opens down.

OpenStudy (dtan5457):

Never did this, but i'll give it a quick shot. give me 5 minutes

OpenStudy (dtan5457):

Got it!

OpenStudy (dtan5457):

your equation of 1/2x^2 start putting your vertex of (-6,2) y coordinate vertex=1/2x^2+2 now plug the -6 for x to get 1/2(x+6)^2+2 simplify the binomial 1/2(x^2+12x+36)+2 halve everything to the parenthesis 1/2x^2+6x+18+2 add the 2 y=1/2x^2+6x+20 plug the vertex of (-6,2) back in, if this equation is right, y=x 2=1/2(-6)^2+6(-6)+20 2=2

OpenStudy (anonymous):

Gotcha! So the equation is y=1/2x^2+6x+20?

OpenStudy (dtan5457):

yes. you can even graph it and it's clearly downwards with a vertex if -6,2

OpenStudy (anonymous):

gotcha! Thanks a ton! Really cleared matters up!

OpenStudy (dtan5457):

I hope so. Honestly, I looked at a question with the answer very similar to this and followed the same instructions for your question.